Version: 3 NASSI

Name

NASSI – Typeset Nassi-Schneiderman diagrams in L^{A}TEX

Date

February 17, 1997

Synopsis

\STRUCT{struct name}{structure purpose}{statements}

For a description of the statements see section Macros

Description

The Nassi macros enable the user to create Nassi-Schneiderman diagrams in a L^{A}TEX document. The macros can be used in any type of document, within all the standard L^{A}TEX environments. An NS-diagram in this context is described in macros resembling a pseudo code. The diagram itself is defined as a structure and is build from statements. Macros are available for all standard programming statements (if, while, action etc.). The Language Sensitive Editor for L^{A}TEX, LSA, is extended to support the creation of NS-diagrams with the Nassi macros. The macros needed to create a Nassi-Schneiderman diagram in your L^{A}TEX document will be made available to you when you include the option Nassi in the \documentstyle command at the start of your L^{A}TEX file, as in the following example:

\documentstyle[11pt,Nassi]{article}

An NS-diagram in the context of the Nassi macros is called a structure. The drawing of the diagram is based on units. A unit can be seen as the box in which an action is described. Both the structure and the unit have an attribute, the width. The layout of the diagram is determined from the value of these attributes. Both can be changed by the user through macros which must be used before the actual structure is described. They have the following effect:

•If both attributes have their default values, the unit width is 60 points (there are 72 points to an inch). The width of the structure is determined by the number of units on one line, while each unit has a minimum width of 60 points. The maximum width of a unit is the structure width.

•If the unit width is specified by the user the same rules apply but the minimum width of a unit is the specified width.

•If the structure width is specified by the user, the diagram will have this width, while the width of each unit is determined depending on their contents and the number of units on one line. The maximum unit width is still the structure width.

It is recommended that users do not change both attributes for the same structure. By default the text in the diagrams is typeset in the same size as the text in the document. This will be to large. Before starting a structure it is adviced to set the desired text size. For the default settings of structure and unit width \scriptsize is recommended.

NASSI Version: 3

When you run into errors like overfull hbox[] you may either have to use a smaller text size like \tiny or give a proper hyphenation directions for the indicated text. Do not forget to reset the text size to normalsize after the structure.

Macros

The available macros fall into four categories:

1.Width settings of structure and unit. They must appear before the structure. 2.Definition of the structure itself. 3.Statements. They must appear within the structure definition. 4.Options. They may appear before as well as within the structure definition.

Width setting macros are:

•\unitwidth=size

Specifies the minimum width of a unit. Size must be given as a standard L^{A}TEX length.

•\nassiwidth=size

Specifies the total width of the diagram. Size must be given as a standard L^{A}TEX length. For a diagram with the same width as the normal text the command \nassiwidth=\textwidth can be used.

Structure definition macro is:

•\STRUCT{struct-name}{struct-description}{statements}%

Specifies the complete diagram. The structure name and description, which can be a short description of its purpose, will be put above the actual diagram. The statements of the structure, described below, must be specified as part of the \STRUCT command.

Statement macros are:

•\ACTION{action}%

Specifies a normal statement.

•\PROC{proc-name}{proc-description}%

Specifies a special kind of action, a procedure call. The proc-name will be put behind the proc-description text in brackets.

•\ACCEPT{entry-name}{statements}\ENDACCEPT%

Specifies the ADA accept statement. The entry-name will be followed by a separate box containing the statements. This statement is not available for flow structures.

•\IF{condition}\THEN{statements}\ELSE{statements}\ENDIF%

Specifies an if statement with a condition and a collection of statements in the THEN and the ELSE part. Neither the THEN nor the ELSE part is optional. If one of them is not needed an empty statement, “{}”, must be used.

•\REPEAT{statements}\UNTIL{test}%

Specifies a loop statement with a collection of statements and an end-condition.

•\WHILE{condition}{statements}\ENDWHILE%

Specifies a loop statement with a start-condition and a collection of statements.

•\CASE{case-item}{when-statements}\ENDCASE%

Specifies a case statement with a case-item and a collection of when-statements. The case statement can also be used to represent the ADA select statement by specifying select as the case-item.

•\WHEN{condition}{statements}%

Part of a case statement. Specifies a condition and a collection of statements. When the case statement is used as an ADA select statement the condition represents the gard, delay or else part.

Options are:

•\setiftext{left-hand}{right-hand}%

Specifies the text to be put on the left-hand and right-hand sides of the condition of an if statement. Defaults are “Y” and “N” for the left-hand and the right-hand respectively. The scope for these settings depends on the position of the command within the document.

–If the command occurs before a \STRUCT statement, the settings remain valid for all \IF statements in all following structures until they are reset by a \setiftext command with the same scope. (Remember to put the \setiftext command after any font size changing commands). –If the command occurs inside a \STRUCT statement but outside any other commands, the settings will remain valid for all \IF statements in the same structure until they are reset by a \setiftext command with the same scope. –If the command occurs inside a command within a structure, the settings will remain valid for all \IF statements within the command (for instance an if statement within the then or else part of another if statement) until they are reset by a \setiftext command with the same scope.

It is allowed to ‘nest’ the \setiftext commands, thus changing the settings for a particular part of a structure or a document.

•\underlinewhentrue% and \underlinewhenfalse%

Specifies whether a line is drawn underneath an \IF, \WHILE, \REPEAT, \ACCEPT or a nested \CASE statement. To be compatible with previous releases of Nassi the default setting is \underlinewhenfalse%. The scope for this settings depends on the position of the command within the document. See \setiftext for more information. This option is not available for flow structures.

Note that each line within the structure, including the command itself, must be concluded by a %. The text appearing (between the braces) in the structure definition and in all statements, including the \setiftext macro is free but must adhere to the standard L^{A}TEX syntax. Every text is considered a paragraph in the L^{A}TEX context.

Files

The file TEX$INPUTS:NASSI.STY specifies the Nassi-Schneiderman macros.

See also

The pages on L^{A}TEX, LSA and FLOW in this section of the manual.

Bugs

•The Nassi option is mutually exclusive with the Flow option because the same macro names are used. •Nassi has a limit to the number of statement macro’s in one structure definition. For example a maximum of 26 IF macro’s can be used. If one of these limits is reached, L^{A}TEX shows some curious error messages, like “You cannot use = as a prefix”, which have no relation at all to the real problem. •Within the \ACCEPT statement part at least one statement is required. As dummy the \ACTION{null} is recommended. •If \underlinewhentrue% is specified and the concerning \WHEN statement is the tallest of all \WHEN statements in the \CASE statement a double line will show. By specifying \underlinewhenfalse% at the end of this \WHEN statement, this problem can be prevented.

Example

The following is an example of a structure with all possible statements. The resulting diagram is shown below.

\scriptsize
\STRUCT{structure name}{structure purpose}{%
\ACTION{initial statement}%
\PROC{proc name}{procedure purpose}%
\ACCEPT{entry name}{%
  \ACTION{critical part}%
}%
\ENDACCEPT%
\IF{condition to test}%
\THEN{%
  \ACTION{true action 1}%
  \ACTION{true action 2}%
}%
\ELSE{%
  \ACTION{false action}%
}%
\ENDIF%
\REPEAT{%
  \ACTION{statement to repeat}%
}%
\UNTIL{end condition}%
\WHILE{start condition}{%
  \ACTION{statement to do}%
}%
\ENDWHILE%
\CASE{case item}{%
  \WHEN{condition 1}{%
    \ACTION{statement to do}%
  }%
  \WHEN{condition 2}{%
    \ACTION{statement 1 to do}%

\ACTION{statement 2 to do}%
}%
\WHEN{condition 3}{%
\ACTION{statement to do}%
}%
}%
\ENDCASE%
\CASE{select}{%
\WHEN{gard 1}{%
\ACCEPT{entry name 1}{%
  \ACTION{statement to do}%
}%
\ENDACCEPT%
}%
\WHEN{}{%
\ACCEPT{entry name 2}{%
  \ACTION{critical statement to do}%
}%
\ENDACCEPT%
\ACTION{non critical statement to do}%
}%
\WHEN{}{%
\ACTION{delay 10.0}%
}%
\WHEN{else}{%
\ACTION{else actions}%
}%
}%
\ENDCASE%
}%
\normalsize

structure name — structure purpose

initial statement procedure purpose (proc name) entry name critical part

[figure: Y; condition to test; N]

true action 1 true action 2

false action

statement to repeat end condition start condition

statement to do

[figure: case item]

condition 1 statement to do condition 2 statement 1 to do statement 2 to do condition 3 statement to do

[figure: select]

gard 1 entry name 1

statement to do entry name 2

critical statement to do non critical statement to do delay 10.0 else else actions